Operation Summary for December 7, 2009 - December 9, 2009
Monday Dec. 7
The day shift began with Operations monitoring the Tevatron (TeV) store 7420 (248.62E30) and the Recycler stash (73.3E10), with the Antiproton Source stacking (14.4E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.
During the day shift, Operators prepared the Meson area enclosures for an access to investigate a ground-faulted magnet in the MT5E string of magnets. The bad magnet was identified. Linac personnel let the I- Source vacuum up to atmosphere and removed the source for repairs. A problem with the Linac Klystron Debuncher water skid caused some beam downtime. MI personnel conducted multibatch slipstacking studies.
During the evening shift, Operations terminated store 7420 at 10:37 PM. Shot setup followed and continued through to the end of the shift.
Tuesday Dec. 8
The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (14.9E10), with Pbar stacking (51.7E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations established store 7421 at 12:23 AM with an average initial luminosity (AIL) of 280E30. Operators discovered an ICW leak above the Booster gradient magnet power supply (GMPS) cabinet in the East Gallery. Operations turned off the Booster beam at 7:30 AM to prepare for repairs.
During the day shift, the Booster ICW leak repair lasted until 1:30 PM. Then, a Booster kicker (B:MP03) developed a current regulation problem, which lasted until the end of the shift.
During the evening shift, A Linac RF station (LRF2) developed problems that held off beam until 10 PM. Operations terminated store 7421 at 10 PM and then put the TeV into shot setup. Operations established store 7422 at 11:55 PM with an AIL of 223.67E30. Antiprotons were stacked (when possible) and transferred to the Recycler. Operations delivered beam to NUMI and MiniBooNE when beam was available.
Wednesday Dec. 9
The midnight shift began with Operations monitoring store 7422 (196.26E30) and the stash (52.7E10), with Pbar stacking (10.4E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations reported numerous oxygen monitor alarms in the TeV due to the low-pressure weather system passing over the Midwest. Operators tuned Booster for better beam positions and to increase intensity.
The Plans for Wednesday
The plans for today are to stack, stash, and store. The Run Coordinator will allow MI, Pbar, and Recycler personnel to conduct studies. TeV personnel will conduct an end-of-store study. The process of rigging in a new magnet for the Meson MTest beamline will begin today. Hopefully, a new MTest experiment will begin taking beam this weekend.
Comments and Suggestions
|last modified 12/09/2009 email Fermilab|
|Security, Privacy, Legal|